OpenStudy (kaylala):

proving identities: (cot+cos)/(sec-cos) = (csc + 1) / tan^2

OpenStudy (solomonzelman):

(cot x+cos x)/(sec x -cos x) = (csc x + 1) / tan^2 ( [cos x / sin x] + cos x ) / ( [1 /cos x] - cos x ) = (csc x + 1) / tan^2 ( [cos x / sin x] + cos x ) / ( [1 /cos x] - cos^2x /cos x ) = (csc x + 1) / tan^2 ------------ ^ I I so far so good?

OpenStudy (kaylala):

wait. i'm lost. i dont get it @SolomonZelman

OpenStudy (solomonzelman):

Alright, do you get the second line that I wrote?

OpenStudy (solomonzelman):

(cot x+cos x)/(sec x -cos x) = (csc x + 1) / tan^2 <--- The initial problem. Translating the left side into cosines and sines ----> ( [cos x / sin x] + cos x ) / ( [1 /cos x] - cos x ) = (csc x + 1) / tan^2 Are you good with this so far?

OpenStudy (kaylala):

i dont get the 2nd line especially the part where the arrow is pointed at

OpenStudy (solomonzelman):

So my first step was translating the left hand side into sines and cosines, where I got the following equation below. ( [cos x / sin x] + cos x ) / ( [1 /cos x] - cos x ) = (csc x + 1) / tan^2

OpenStudy (solomonzelman):

then the arrow pointing is just to attract your attention to where am I changing the equation, I am making cos x into cos^2x/cosx to add inside the second parenthesis.

OpenStudy (kaylala):

okay got it now @SolomonZelman then what?

OpenStudy (anonymous):